OpenStudy (anonymous):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Make 6 squares. The squares are numbered sequentially from 1 to 6. Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing an even tile from the first bag and an even tile from the second bag? Answers are 6/12 9/12 6/36 and 9/36 I was think it was 6/12 am I right?

OpenStudy (anonymous):

I think its 6/12

OpenStudy (anonymous):

but no sure

OpenStudy (anonymous):

he had 3/6 chance of drawing even from each bag. Multiply the results 3/6 * 3/6 =

OpenStudy (anonymous):

6/36

OpenStudy (anonymous):

\[\frac{ 3 }{ 6 } \times \frac{ 3 }{ 6 }=\frac{ 3\times3 }{ 6\times6 }=\frac{ 9 }{ }\]

OpenStudy (anonymous):

9/36

OpenStudy (anonymous):

wait my bad I ment 9

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome